Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
A rectangular blanket is 20 inches longer than it is wide. If the area of the blanket is 1,125 square inches, find the length and the width. Show work
i think this one is l*w*h but some one please tell me for sure

Answer by mananth(16946)

Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
A rectangular blanket is 20 inches longer than it is wide. If the area of the blanket is 1,125 square inches, find the length and the width. Show work
i think this one is l*w*h but some one please tell me for sure
Let the width be x
the length will be x+20
Area of rectangle = L*W
x*(x+20)= 1125
x^2+20x=1125
x^2+20x-1125=0
x^2+45x-25x-1125=0
x(x+25)-25(x+45)=0
(x-25)(x+45)=0
x= 25 OR -45
the width cant be -ve value
so width = 25 inches
Length will be 25+20 =45 Inches